Difference in final temperatures? Please show all steps

Two containers (call them A and B) with the same initial volume, V 31 L, hold the same amount of monatomic ideal gas, say helium, at the same initial temperature T = 301 K and the same initial pressure Pi 1 atm. Container A holds the gas at a constant volume and Container B holds the gas at a constant pressure If you add Q 5389 J of heat to both containers, what is the difference in the final temperatures after the heat is added, TfA TB? This is the difference in the final temperatures between Container A, a constant Volume process, and Container B, a constant Pressure process, when the heat added is the same Give your answer in Kelvin to three significant digits. The amount of gas in the containers does not change.
Answer #1

DATA:

\\ V_{i,A}=V_{i.B}=31\,L \\ T_{i,A}=T_{i,B}=301 K \\ P_{i,A}=P_{i,B}=1 \,atm \\ Q=5389\,J

Solution:

According to the gas ideal equation:

PV=nRT\: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: (1)

Where:

P: is the Pressure in the containers, which are equal.

V: is the volume

n: is the number of mole

R: is the ideal gas constant, which has a value of:

R=0.0821\,\frac{atm.L}{mol.K}\: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: (2)

T: is the temperature.

Now, isolating n from equation (1) we have:

n=\frac{PV}{RT}\: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: (3)

The number of mole in both containers are the same, so:

n_{A}=n_{B}=\frac{PV}{RT}\: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: (4)

Inserting (2) and data given values into equation (4) we have:

n_{A}=n_{B}=\frac{(1\,atm)(31\,L)}{(0.0821\,\frac{atm.L}{mol.K})(301\,K)}\: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \:

Solving we obtain:

n_{A}=n_{B}=1.26\,mol\: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: (5)

Now, the heat added on each container is defined as:

  • For container A . Constant Volume we have:

Q=n_{A}C_{V}\Delta T

or

Q=n_{A}C_{V}\left ( T_{f,A}-T_{i,A} \right )

Isolating T_{f,A} from equation above we have:

T_{f,A}=T_{i,A}+\frac{Q}{n_{A}C_{V}}\: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: (6)

Where, C_{V} is the molar specific heat at a constant volume, which has a value of:

C_{V}=\frac{3}{2}R=12.5\,\frac{J}{mol.K}\: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: (7)

Now, replacing (5), (7) and data given values into equation (6) we get:

T_{f,A}=301\,K+\frac{5389\,J}{(1.26\,mol)(12.5\, \frac{J}{mol.K})}\: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \:

Solving we obtain:

{\color{Red} T_{f,A}=643\,K}\: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: (8)

  • For Container B, Constant Presurre we have:

C_{P}=\frac{5}{2}R=20.8\,\frac{J}{mol.K}\: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: (9)

Therefore, the final temperature on the container B is:

T_{f,B}=T_{i,B}+\frac{Q}{n_{B}C_{P}}\: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: (10)

Replacing (5) , (9) and data given values into equation (10) we have:

T_{f,B}=301\,K+\frac{5389\,J}{(1.26\,mol)(20.8\,J)}\: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: (11)

Solving we obtain:

{\color{Red} T_{f,B}=507\,K}\: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: \: (12)

Now, the difference in the final temperatures is :

T_{f,A}-T_{f,B}=643\,K-507\,K

Solving we obtain:

{\color{Blue} T_{f,A}-T_{f,B}=136\,K}